Sales price $6.5 per unit Variable costs $2.64 per unit Fixed Costs $10,816 Budgeted number of units 4,648 What is margin of safety in units?
Answer #1

Margin of safety = 1,846 Units

Working

Margin of safety formula = Actual sales units - Breakeven sales units.

A Sales price $               6.50
B Variable cost per unit $               2.64
C=A-B Contribution per unit $               3.86
D Fixed cost $    10,816.00
E=D/C Break-even Sales in units                 2,802
F Actual units                 4,648
G=F-G Margin of safety                 1,846